OpenStudy (anonymous):

Use the Laplace transform to solve y^(4)-4y'''+6y"-4y'+y=0; y(0)=0, y'(0)=1, y"(0)=0, y'''(0)=1. ( I got L(y^(4))-4L(y''')+6L(y'')-4L(y')+L(y)=L(0) but I don't know what's L(y^(4)) and L(y'''), please help.)

hartnn (hartnn):

\(\huge \color {green }{L(y^n)=s^nL(y)-\sum \limits_{k=1}^ns^{k-1}y^{n-k}(0)}\)

hartnn (hartnn):

L (y') = sL(y) -y(0) L(y'')= s^2 L(y) -s y(0) -y'(0) L(y''') = s^3 L(y) -s^2 y(0) -s y'(0) - y''(0) L(y^4) = s^4 L(y) -s^3 y(0) -s^2 y'(0)-s y''(0)-y'''(0)

hartnn (hartnn):

phew....

OpenStudy (anonymous):

Wait a minute, please don't leave.

OpenStudy (anonymous):

Thanks for being patient, please wait a little more. I'm almost done.

hartnn (hartnn):

ok, do you need me to verify your answer ? because i wasn't solving in the mean time :P

OpenStudy (anonymous):

It's alright, let me finish the problem.

OpenStudy (anonymous):

Okay, so I got (s^2+4s-5)/(s^4-4s^3+6s^2-4s+1), what to do next?

hartnn (hartnn):

try to factorize s^4-4s^3+6s^2-4s+1 if you remember pascal triangle, you will be able to do it in a blink of an eye

OpenStudy (anonymous):

But how to do the pascal triangle for this one?

hartnn (hartnn):

the co-efficients are 1,-4,6,-4,1 which would make that \( \large (s+(-1))^4=(s-1)^4 \)

OpenStudy (anonymous):

Could you do synthetic division?

hartnn (hartnn):

after guessing one root as +1, yes. finally you'll get (x-1)^4

OpenStudy (anonymous):

Wait like 3 minutes, I'll be back.

OpenStudy (anonymous):

So you get (s+5)/(s-1)^3, right?

hartnn (hartnn):

yup, correct :) know what to do next ?

OpenStudy (anonymous):

To take the laplace transform? I need to look at the table, right?

hartnn (hartnn):

yeah, but you need to apply one property first...

OpenStudy (anonymous):

What is it?

hartnn (hartnn):

ever seen something like this? \(L^{-1}(F(s+a))=e^{-at}L^{-1}F(s)\) ?

OpenStudy (anonymous):

No. But how does that apply to the problem?

hartnn (hartnn):

you have s-1 in the denominator, so e^x will come out \(\large e^x L^{-1}[\dfrac{(s+1)+5}{s^3}]\)

OpenStudy (anonymous):

Sorry, I don't understand what you wrote.

hartnn (hartnn):

then how were you thinking og getting inverse laplace transform of (s+5)/(s-1)^3 ?

OpenStudy (anonymous):

I have no idea.

hartnn (hartnn):

thats why i gave you the property, which needs to be used here.... \(L^{-1}(F(s+a))=e^{-at}L^{-1}F(s)\)

hartnn (hartnn):

to make , s+a as just s, we have to multiply e^-at or, s-a would require e^at here, to have just s^3 instead of (s-1)^3 , we will multiply e^x (e^(+1x)) to inverse laplace after replacing s by s+1

OpenStudy (anonymous):

I got it. Thanks for your time.

hartnn (hartnn):

oh, good! welcome ^_^

OpenStudy (anonymous):

You're right. Thanks, my friend. Can you help me with other question? http://openstudy.com/study#/updates/52056454e4b01dab33705ac7

OpenStudy (loser66):

@hartnn Please , explain more about them. I would like to know. \[L^{-1}(F(s+a))=e^{-at}L^{-1}F(s)\]\(\rightarrow\)\[\large e^x L^{-1}[\dfrac{(s+1)+5}{s^3}]\]